# Equality of Hölder exponents for distribution functions of Gibbs measures

### v1

@@ 3. The main theorem
Suppose that $P(\psi)=0$. Suppose that there exists a finite set $A\subset\mathbb{R}$ and $\ell_{0}\geq 1$ such that for all $\ell\geq\ell_{0}$ and all $k_{1},\dots,k_{\ell}\in I$ we have
@@
Assume that $\mathcal{P}(\psi)=0$, and $\psi$ is not cohomologous to $\varphi$. Then, for every $k\in\mathbb{N}$ and $x\in\mathbb{R}$ the limit
@@
Let $\omega\in I^{\mathbb{N}}$ such that $\pi(\omega)=x$. Let $s_{n},t_{n}]:=\pi\left(\left[\omega_{n}\right]\right)$. Then $s_{n}\leq x\leq t_{n}$, and
@@
Indeed such $\tau_{1}\dots\tau_{\ell}$ always exist for otherwise, Lemma 3.3 implies that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $\dim_{H}\Lambda=k=1$, contradicting our hypothesis.
@@
Recall that the maps $\phi_{i},i\in I$ are all increasing, so if $u$ and $v$ are two words of the same length and $u\prec v$ in the lexicographical order, then $\pi([u])$ lies to the left of $\pi([v])$.
We claim that there is an infinite subset $\mathcal{N}\subset\mathbb{N}$ such that for each $n\in\mathcal{N}$, the point $x$ and the interval $s_{n}^{N},t_{n}^{N}]$ are separated by at least a full cylinder interval of level at most $n+\ell+1$. To see this, let $j<\ell$ be the smallest integer such that $\tau_{j+1}\neq\tau_{1}$. Such $j$ exists since $\tau_{1}\dots\tau_{\ell}$ contains at least two different letters. Write $\tau_{1}\dots\tau_{\ell}=\tau_{1}^{j}\tau_{j+1}\dots\tau_{\ell}$, and consider two cases:
Case 1. If $\omega$ ends in $\tau_{1}^{\infty}$, then for every sufficiently large $n\in\mathbb{N}$ the cylinder interval $\pi\big{(}[\omega|_{n}\tau_{1}^{j+1}\tau_{j+1}]\big{)}$ separates $x$ from the interval $\pi\big{(}[\omega|_{n}\tau_{1}\dots\tau_{j+1}]\big{)}$, and hence from $s_{n}^{N},t_{n}^{N}]$.
Case 2. Otherwise, there are infinitely many $n$ such that $\omega_{n+1}\neq\tau_{1}$. Take such an $n$. If $\omega_{n+1}<\tau_{1}$, then
$\omega_{n+1}\dots\omega_{n+\ell+1}\prec\tau_{1}1^{\ell}\prec\tau_{1}\dots\tau_{\ell}\tau_{1},$
so the cylinder interval $\pi\big{(}[\omega|_{n}\tau_{1}1^{\ell}]\big{)}$ lies between $x$ and $s_{n}^{N},t_{n}^{N}]$. A similar argument, with the inequalities reversed, applies when $\omega_{n+1}>\tau_{1}$.
In either case, the claim follows. By replacing $\mathcal{N}$ by an infinite subset if necessary, we may assume that either $t_{n}^{N}>x$ for all $n\in\mathcal{N}$, or $t_{n}^{N}<x$ for all $n\in\mathcal{N}$. Without loss of generality, we assume the former. The argument in the latter case is similar. Thus, $r_{n}^{N}>0$ for all $n\in\mathcal{N}$.
@@
(13) $\mathrm{diam}\pi([uv])\geq\mathrm{diam}\pi([u])\cdot r_{\min}^{|v|}$
@@
For $n\in\mathcal{N}$, let $u_{n},v_{n}]$ be a cylinder interval of length $n+\ell+1$ separating $x$ and $s_{n}^{N},t_{n}^{N}]$. Then by (13),
$|t_{n}^{N}-x|\geq v_{n}-u_{n}\geq\mathrm{diam}\pi([\omega|_{n}])\cdot r_{\min}^{\ell+1}.$
@@
(15) $r_{n}^{N}=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}\asymp_{\ell}\frac{\mathrm{diam}\left(\pi\left(\left[\omega|_{n}\right]\right)\right)}{\mathrm{diam}\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}\asymp_{\ell}e^{-NS_{\ell}\left(\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
@@
On the other hand, by the Gibbs property (6) of $\mu_{\psi}$ and the bounded distortion property (12) we have
$\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}$ $=\frac{\mu_{\psi}\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}{\left(\mathrm{diam}\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)\right)^{k}}$
$\asymp_{\ell}\frac{\mu_{\psi}\left(\left[\omega|_{n}\right]\right)}{\left(\mathrm{diam}\pi\left(\left[\omega|_{n}\right]\right)\right)^{k}}\cdot e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}$
@@
$2^{1-k}L\leq\big{(}(r^{N})^{k}+(1-r^{N})^{k}\big{)}L=\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}\asymp e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
@@ 4. The pointwise Hölder spectrum of $F_{\mu}$
Recall from the Introduction that $E_{F}(\alpha)=\{t\in\mathbb{R}:\alpha_{F}(t)=\alpha\}$. As a consequence of Theorem 3.1, we obtain the following result.
@@
For $\alpha\geq\alpha_{0}$, we trivially have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\geq\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$ since $\mathscr{E}_{*}(\alpha;\mu)\supset\mathscr{E}(\alpha,\mu)$. But we also have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\leq\beta^{*}(\alpha)$. This follows directly from a careful examination of the proof that $\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$; see, for instance, [21, p. 501]. We thus conclude that

### v2

@@ 3. The main theorem
Let $\mathbf{i},\mathbf{j}\in I^{p}$ with $\mathbf{i}\neq\mathbf{j}$, where $p\geq 3$. Let $n:=|\mathbf{i}\wedge\mathbf{j}|+1$ be the first coordinate at which $\mathbf{i}$ and $\mathbf{j}$ disagree. Assume $n\leq p-2$, and suppose the cylinder intervals $\pi([\mathbf{i}])$ and $\pi([\mathbf{j}])$ have a common endpoint. Then $|i_{n}-j_{n}|=1$ and one of the following holds:
  1. (i)
    $i_{n+1}\dots i_{p}\in\big\{1^{p-n},m^{p-n},1m^{p-n-1},m1^{p-n-1}\big\}$; or
  2. (ii)
    $i_{n+1}\dots i_{p}$ is a prefix of $(1m)^{\infty}$ and $i_{n}\neq m$; or
  3. (iii)
    $i_{n+1}\dots i_{p}$ is a prefix of $(m1)^{\infty}$ and $i_{n}\neq 1$.
Furthermore, if $\mathbf{i}$ satisfies (ii) or (iii), then so does $\mathbf{j}$.
Proof.
Since the cylinder intervals $\pi([1]),\dots,\pi([m])$ are arranged from left to right, the cylinder intervals $\pi([\mathbf{i}|_{n-1}1]),\dots,\pi([\mathbf{i}|_{n-1}m])$ are arranged either from left to right or from right to left in $\pi([\mathbf{i}|_{n-1}])$ (depending on the orientation of $\phi_{i_{1}}\circ\dots\circ\phi_{i_{n-1}}$). Hence, $|i_{n}-j_{n}|=1$.
Assume without loss of generality that $\pi([\mathbf{i}|_{n}])$ lies to the left of $\pi([\mathbf{j}|_{n}])$. Then for each $l=n+1,\dots,p$, $\pi([\mathbf{i}|_{l}])$ is the rightmost subcylinder of $\pi([\mathbf{i}|_{l-1}])$ and $\pi([\mathbf{j}|_{l}])$ is the leftmost subcylinder of $\pi([\mathbf{j}|_{l-1}])$. Thus, $i_{l},j_{l}\in\{1,m\}$ for $l=n+1,\dots,p$.
If $\phi_{1}$ is increasing, then in $i_{n+1}\dots i_{p}$ any digit $1$ must be succeeded by another $1$. Similarly, if $\phi_{m}$ is increasing, any digit $m$ must be followed by another $m$. On the other hand, if $\phi_{1}$ is decreasing, any digit $1$ must be followed by a digit $m$; and likewise, if $\phi_{m}$ is decreasing, any digit $m$ must be followed by a digit $1$. This implies that $i_{n+1}\dots i_{p}$ is of one of the forms in (i) or else is a prefix of either $(1m)^{\infty}$ or $(m1)^{\infty}$.
Suppose $i_{n+1}\dots i_{p}$ is a prefix of $(1m)^{\infty}$. Then $\phi_{1}$ and $\phi_{m}$ are both decreasing. If $i_{n}=m$, then necessarily $j_{n}=m-1$. Since $\pi([\mathbf{i}|_{n}])$ lies to the left of $\pi([\mathbf{j}|_{n}])$, this means that the map $\phi_{i_{1}}\circ\dots\circ\phi_{i_{n-1}}$ is decreasing. But $\phi_{m}$ is also decreasing, and so $\phi_{i_{1}}\circ\dots\circ\phi_{i_{n}}$ is increasing. Therefore, the rightmost subcylinder of $\pi([\mathbf{i}|_{n}])$ is $\pi([\mathbf{i}|_{n}m])$, contradicting that $i_{n+1}=1$. Hence, $i_{n}\neq m$. A similar argument shows that if $i_{n+1}\dots i_{n}$ is a prefix of $(m1)^{\infty}$, then $i_{n+1}\neq 1$. ∎
Lemma 3.5.
Let $\ell\geq 2$ and let $\tau_{1}\dots\tau_{\ell}\in I^{\ell}$ be a word such that $\tau_{1}\neq\tau_{2}$. Then for any $\omega\in I^{\mathbb{N}}$ there are infinitely many integers $n$ such that the cylinder intervals $\pi([\omega|_{n+3\ell}])$ and $\pi([\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{3}])$ do not have a common endpoint.
Proof.
Let $\omega\in I^{\mathbb{N}}$, and consider two cases:
Case 1. Suppose first that there exist infinitely many $n$ such that $\omega_{n}\neq\tau_{1}$. Take such an $n$. Suppose, by way of contradiction, that the cylinder intervals $\pi([\omega|_{n-1}\omega_{n}\dots\omega_{n+\ell+1}])$ and $\pi([\omega|_{n-1}\tau_{1}\dots\tau_{\ell}\tau_{1}\tau_{2}])$ are adjacent. Then either (i), (ii) or (iii) of Lemma 3.4 holds for the word $\mathbf{i}:=\omega|_{n-1}\tau_{1}\dots\tau_{\ell}\tau_{1}\tau_{2}$, with $p:=n+\ell+1$, so that $i_{n+1}\dots i_{p}=\tau_{2}\dots\tau_{\ell}\tau_{1}\tau_{2}$. Since $\tau_{2}\neq\tau_{1}$, we can rule out (i). If $\tau_{2}\dots\tau_{n}\tau_{1}\tau_{2}$ is a prefix of $(1m)^{\infty}$, then $\tau_{2}=1$ and so $\tau_{1}=m$, which means (ii) does not hold. By a symmetric argument, (iii) does not hold. Hence, $\pi([\omega|_{n+\ell+1}])$ and $\pi([\omega|_{n-1}\tau_{1}\dots\tau_{\ell}\tau_{1}\tau_{2}])$ are not adjacent, so neither are $\pi([\omega|_{n-1+3\ell}])$ and $\pi([\omega|_{n-1}(\tau_{1}\dots\tau_{\ell})^{3}])$, since $\ell+2\leq 3\ell$.
Case 2. Suppose $\omega$ ends in $\tau_{1}^{\infty}$. Then $\omega_{n-1}=\tau_{1}$ and $\omega_{n}\neq\tau_{2}$ for all large enough $n$. Fix such a large $n$. We have two further cases:
Case 2a. Assume first that $\ell\geq 3$. Apply Lemma 3.4 with $p:=n+\ell+1$, $\mathbf{i}:=\omega|_{n-1}\tau_{2}\dots\tau_{\ell}\tau_{1}\tau_{2}\tau_{3}=\omega|_{n-2}\tau_{1}\dots\tau_{\ell}\tau_{1}\tau_{2}\tau_{3}$ and $\mathbf{j}:=\omega|_{n+\ell+1}$, so $i_{n+1}\dots i_{p}=\tau_{3}\dots\tau_{\ell}\tau_{1}\tau_{2}\tau_{3}$. We can immediately rule out condition (i) of Lemma 3.4, because $i_{n+1}=i_{p}$ and $i_{p-2}=\tau_{1}\neq\tau_{2}=i_{p-1}$. On the other hand, if $\tau_{3}\dots\tau_{n}\tau_{1}\tau_{2}\tau_{3}$ is a prefix of $(1m)^{\infty}$, then $\tau_{3}=1$ and hence $i_{n}=\tau_{2}=m$, so (ii) does not hold. Similarly, (iii) does not hold. Thus, the cylinder intervals $\pi([\mathbf{i}])$ and $\pi([\mathbf{j}])$ are not adjacent, so neither are $\pi([\omega|_{n-2+3\ell}])$ and $\pi([\omega|_{n-2}(\tau_{1}\dots\tau_{\ell})^{3}])$, since $\ell+3\leq 3\ell$.
Case 2b. Finally, assume $\ell=2$. We then replace the word $\tau_{2}\dots\tau_{\ell}\tau_{1}\tau_{2}\tau_{3}$ in Case 2a with $\tau_{2}\tau_{1}\tau_{2}\tau_{1}$, and argue similarly. ∎
Lemma 3.6.
Suppose that there exists a finite set $A\subset\mathbb{R}$ and $\ell_{0}\geq 1$ such that for all $\ell\geq\ell_{0}$ and all $k_{1},\dots,k_{\ell}\in I$ we have
@@
Since $A$ is finite, we have $\alpha_{-}=\alpha_{+}$. By a well-known dichotomy for multifractal spectra [21], we have that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$. Clearly, this implies $A=\left\{\dim_{H}\Lambda\right\}$. ∎
@@
For every $k\in\mathbb{N}$ and $x\in\mathbb{R}$ the limit
@@
Let $\omega\in I^{\mathbb{N}}$ such that $\pi(\omega)=x$. Let $s_{n},t_{n}]:=\pi\left(\left[\omega|_{n}\right]\right)$. Then $s_{n}\leq x\leq t_{n}$, and
@@
Indeed such $\tau_{1}\dots\tau_{\ell}$ always exist for otherwise, Lemma 3.6 implies that $\psi$ is cohomologous to $(\dim_{H}\Lambda)\varphi$ and $\dim_{H}\Lambda=k=1$, contradicting our hypothesis. Replacing $\tau_{1}\dots\tau_{\ell}$ by a cyclical permutation if necessary, we may assume that $\tau_{2}\neq\tau_{1}$.
@@
By Lemma 3.5 there is an infinite subset $\mathcal{N}\subset\mathbb{N}$ such that for each $n\in\mathcal{N}$, the point $x$ and the interval $s_{n}^{N},t_{n}^{N}]$ are separated by a full cylinder interval of level $n+3\ell$. Replacing $\mathcal{N}$ by an infinite subset if necessary, we may assume that either $t_{n}^{N}>x$ for all $n\in\mathcal{N}$, or $t_{n}^{N}<x$ for all $n\in\mathcal{N}$. Without loss of generality, we assume the former. The argument in the latter case is similar. Thus, $r_{n}^{N}>0$ for all $n\in\mathcal{N}$.
@@
(11) $\mathrm{diam}(\pi([uv]))\geq\mathrm{diam}(\pi([u]))\cdot r_{\min}^{|v|}$
@@
For $n\in\mathcal{N}$, let $u_{n},v_{n}]$ be a cylinder interval of length $n+3\ell$ separating $x$ and $s_{n}^{N},t_{n}^{N}]$. Then by (11),
$|t_{n}^{N}-x|\geq v_{n}-u_{n}\geq\mathrm{diam}(\pi([\omega|_{n}]))\cdot r_{\min}^{3\ell}.$
@@
(13) $r_{n}^{N}=\frac{t_{n}^{N}-x}{t_{n}^{N}-s_{n}^{N}}\asymp_{\ell}\frac{\mathrm{diam}\left(\pi\left(\left[\omega|_{n}\right]\right)\right)}{\mathrm{diam}(\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}\asymp_{\ell}e^{-NS_{\ell}\left(\varphi\right))(\overline{\tau_{1}\dots\tau_{\ell}})}.$
@@
On the other hand, by the Gibbs property (5) of $\mu_{\psi}$, the bounded distortion property (4) and (10) we have
$\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}$ $=\frac{\mu_{\psi}\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right]\right)}{\left(\mathrm{diam}(\pi\left(\left[\omega|_{n}(\tau_{1}\dots\tau_{\ell})^{N}\right])\right)\right)^{k}}$
$\asymp_{\ell}\frac{\mu_{\psi}\left(\left[\omega|_{n}\right]\right)}{\left(\mathrm{diam}(\pi\left(\left[\omega|_{n}\right])\right)\right)^{k}}\cdot e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}$
@@
$2^{1-k}L\leq\big((r^{N})^{k}+(1-r^{N})^{k}\big)L=\lim_{n\to\infty}\frac{F(t_{n}^{N})-F(s_{n}^{N})}{(t_{n}^{N}-s_{n}^{N})^{k}}\asymp_{\ell}e^{NS_{\ell}\left(\psi-k\varphi\right)(\overline{\tau_{1}\dots\tau_{\ell}})}.$
@@ 4. The pointwise Hölder spectrum of $F_{\mu}$
Recall from the Introduction that $E_{F}(\alpha)=\{t\in\mathbb{R}:\alpha_{F}(t)=\alpha\}$. As a consequence of Theorem 3.1, we obtain the following result, in which $\alpha_{-}$ and $\alpha_{+}$ are given by (6), and $\beta^{*}(\alpha)$ is defined by (8).
@@
For $\alpha\geq\alpha_{0}$, we trivially have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\geq\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$ since $\mathscr{E}_{*}(\alpha;\mu)\supset\mathscr{E}(\alpha,\mu)$. But we also have $\dim_{P}\mathscr{E}_{*}(\alpha;\mu)\leq\beta^{*}(\alpha)$. This follows directly from a careful examination of the proof that $\dim_{P}\mathscr{E}(\alpha;\mu)=\beta^{*}(\alpha)$; see, for instance, [19, p. 501]. We thus conclude that
@@
5. Conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ maps
We apply our results to conjugacy maps between expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval maps, extending [9, Corollary 2.3] in two ways. Firstly, we also analyze the pointwise Hölder exponent $\alpha$ of the conjugacy map, whereas in [9] only $\tilde{\alpha}$ was considered. Secondly, we allow interval maps with orientation-reversing branches.
Let $f$ be an expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval map with $s\geq 2$ full branches, i.e., there exist closed intervals $J_{1},\dots,J_{s}\subset[0,1]$ with non-empty, pairwise disjoint interiors such that $f|_{J_{i}}$ has a $\mathcal{C}^{1+\epsilon}$ extension to a neighbourhood of $J_{i}$ satisfying $|f^{\prime}|_{J_{i}}|>1$ and $f(J_{i})=[0,1]$, for $1\leq i\leq s$. We will always assume that $\sup J_{i}\leq\inf J_{j}$ if $i<j$. The repeller of $f$ is denoted by $\Lambda$ and the restriction $f|_{\Lambda}:\Lambda\rightarrow\Lambda$ is semi conjugate to the shift $\sigma:I^{\mathbb{N}}\rightarrow I^{\mathbb{N}}$ with alphabet $I:=\left\{1,\dots,s\right\}$. The semi conjugacy is given by the coding map $\pi_{f}:I^{\mathbb{N}}\rightarrow\Lambda$ of the conformal iterated function systems $\Phi_{f}$, which is given by the contracting inverse branches $(f|_{J_{i}})^{-1}:[0,1]\rightarrow[0,1]$, $1\leq i\leq s$. Note that each branch is either orientation-preserving, or orientation-reversing. We denote by $\varphi_{f}:I^{\mathbb{N}}\rightarrow\mathbb{R}$ the geometric potential of $\Phi_{f}$ given by $\varphi_{f}:=-\log|f^{\prime}\circ\pi_{f}|$.
We say that two piecewise $\mathcal{C}^{1+\epsilon}$ interval maps $f$ and $g$ have the same signature if they share the same alphabet $I$ and satisfy
$\{i\in I\mid f_{i}\text{ is orientation-preserving }\}=\{i\in I\mid g_{i}\text{ is orientation-preserving }\}.$
If $f$ and $g$ have the same signature and $\Lambda_{g}=[0,1]$, then $\pi_{g}(\pi_{f}^{-1}\{x\})$ is a singleton for each $x\in\Lambda_{f}$. We can thus define the conjugacy map
$\Theta:\Lambda_{f}\rightarrow[0,1],\quad\{\Theta(x)\}=\pi_{g}(\pi_{f}^{-1}\{x\}).$
Slightly abusing notation, we write $\Theta=\pi_{g}\circ\pi_{f}^{-1}$. Note that we have $\Theta\circ f=g\circ\Theta$, which implies for each $x\in\Lambda_{f}$,
(18) $\pi_{g}^{-1}\left((-\infty,\Theta(x)]\right)=\pi_{f}^{-1}\left((-\infty,x]\right).$
The conjugacy $\Theta$ has a unique continuous extension to a function on $\mathbb{R}$ which is locally constant on $\mathbb{R}\setminus\Lambda_{f}$. We denote this function again by $\Theta$ and refer to it as the conjugacy map between $f$ and $g$. Note that (18) continues to hold for the conjugacy map between $f$ and $g$.
Since $\Lambda_{g}=[0,1]$ we have that the Lebesgue measure $\lambda$ on $[0,1]$ satisfies $\lambda=\mu_{\varphi_{g}}\circ\pi_{g}^{-1}$ where $\mu_{\varphi_{g}}$ is a Gibbs measure for $\varphi_{g}$. Hence, by (18), the conjugacy map $\Theta$ between $f$ and $g$ coincides with the distribution function of $\mu_{\varphi_{g}}\circ\pi_{f}^{-1}$. Our main theorem thus implies the following.
Corollary 5.1.
Let $f$ and $g$ be two expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval maps with full branches and with the same signature. Suppose that $\varphi_{f}$ is not cohomologous to $\varphi_{g}$. Then, for the conjugacy map $\Theta:\Lambda_{f}\rightarrow[0,1]$ between between $f$ and $g$ we have
$\dim_{H}\left\{t\in\Lambda\mid\alpha_{\Theta}(t)=\alpha\right\}=\dim_{H}\left\{t\in\Lambda\mid\tilde{\alpha}_{\Theta}(t)=\alpha\right\}=\beta^{*}(\alpha),\quad\alpha\in\left[\alpha_{-},\alpha_{+}\right],$
where $\beta:\mathbb{R}\rightarrow\mathbb{R}$ is defined implicitly by $\mathcal{P}(\beta(q)\varphi_{f}+q\varphi_{g})=0$, and $\alpha_{-}<\alpha_{+}$.
Remark 5.2.
Let $f$ and $g$ be two expanding piecewise $\mathcal{C}^{1+\epsilon}$ interval maps with full branches. Suppose in addition that all branches of $f$ and $g$ are monotonically increasing and that $\Lambda_{f}=\Lambda_{g}=[0,1]$. Then, it follows from [13, Proof of Theorem 1.2] that the conjugacy map $\Theta:[0,1]\rightarrow[0,1]$ between $f$ and $g$ is a $\mathcal{C}^{1+\epsilon}$-diffeomorphism if and only if $\varphi_{f}$ is cohomologous to $\varphi_{g}$.
Example 5.3.
Let
$f(x):=\begin{cases}3x&\mbox{if $0\leq x\leq 1/3$},\\ 3x-2&\mbox{if $2/3\leq x\leq 1$},\end{cases}\qquad g(x):=\begin{cases}2x&\mbox{if $0\leq x<1/2$},\\ 2x-1&\mbox{if $1/2\leq x\leq 1$}.\end{cases}$
Then the repeller $\Lambda$ of $f$ is the ternary Cantor set and the conjugacy $\Theta$ between $f$ and $g$ is the ternary Cantor function. Suppose we invert the second branch of both $f$ and $g$; that is, we consider instead the functions
$\tilde{f}(x):=\begin{cases}3x&\mbox{if $0\leq x\leq 1/3$},\\ 4-3x&\mbox{if $2/3\leq x\leq 1$},\end{cases}\qquad\tilde{g}(x):=\begin{cases}2x&\mbox{if $0\leq x<1/2$},\\ 3-2x&\mbox{if $1/2\leq x\leq 1$}.\end{cases}$
Note that $\tilde{f}$ and $\tilde{g}$ again have the same signature. The repeller of $\tilde{f}$ is still the ternary Cantor set, and the conjugacy $\tilde{\Theta}$ between $\tilde{f}$ and $\tilde{g}$ is still the ternary Cantor function. In other words, $\tilde{\Theta}:=\pi_{\tilde{g}}\circ\pi_{\tilde{f}}^{-1}=\Theta$. This follows because, if the coding of $x\in\Lambda$ under $f$ is $\omega:=(\omega_{n})_{n}:=\pi_{f}^{-1}(x)$, then the coding of $x$ under $\tilde{f}$ is $\tilde{\omega}=(\tilde{\omega}_{n})_{n}=\pi_{\tilde{f}}^{-1}(x)$, where $\tilde{\omega}_{1}=\omega_{1}$ and successively, for $n=2,3,\dots$,
$\tilde{\omega}_{n}:=\begin{cases}\omega_{n}&\mbox{if $\#\{i<n:\tilde{\omega}_{i}=2\}$ is even},\\ 3-\omega_{n}&\mbox{otherwise},\end{cases}$
[…]
Example 5.4.
More generally, let $m,n\in\mathbb{N}$ with $m,n\geq 3$. Let
$f(x):=\begin{cases}mx&\mbox{if $0\leq x\leq 1/m$},\\ nx-(n-1)&\mbox{if $(n-1)/n\leq x\leq 1$},\end{cases}$
and let $g(x)$ again be as in Example 5.3. If $m=n$, the situation is much as it was in the last example. But if $m\neq n$, the conjugacy $\Theta$ changes when we reverse the second branch of both $f$ and $g$, because the repeller $\Lambda$ changes. Nonetheless, the new conjugacy $\tilde{\Theta}$ has the same Hölder spectrum as $\Theta$. The range of the spectrum is
$\alpha_{-},\alpha_{+}]=\left[\min\left\{\frac{\log 2}{\log m},\frac{\log 2}{\log n}\right\},\max\left\{\frac{\log 2}{\log m},\frac{\log 2}{\log n}\right\}\right],$
and the spectrum is given by Corollary 5.1.
Acknowledgments
P.A. was partially supported by Simons Foundation grant #709869. J.J. was partially supported by the JSPS KAKENHI 24K06777.